Calculate work done by a man of mass 60kg who climbs up a stair consisting of 15 step.Given of each steps is 30cm.


1
Expert's answer
2020-11-05T10:07:31-0500

The weight of the man (and, thus, the foce he applies against the gravity) is:


"P = mg = 60kg\cdot 9.81m/s^2 = 588.6N"

The work done by this force while climbing on one step is:


"A_1 = Ph = 588.6N\times 0.3m = 176.58\space J"

where "h = 30cm = 0.3m" is a height of the step. The work done on the whole stair is:


"A = 15\times A_1 = 2648.7\space J"

Answer. 2648.7 J